OpenStudy (anonymous):

−2(x + 3) = −4(x − 6)

OpenStudy (anonymous):

get rid of the brackets

OpenStudy (anonymous):

And after that?

OpenStudy (anonymous):

-2x-6 = -4x+24 -2x+4x=24+6 2x=30 x=15

OpenStudy (anonymous):

Thank you